OpenStudy (anonymous):

HELP! A person earns $16,700 one year and gets a 5% raise in salary. What is the new salary?

OpenStudy (anonymous):

16700*105/100

OpenStudy (anonymous):

or salary is 16700 and raise is 16700*(5/100) Total = salary + raise = 16700 + 16700(5/100) = 16700*(105/100)

OpenStudy (anonymous):

Thank you!!!